# In-depth characterization of the Kepler-10 three-planet system with HARPS-N RVs and Kepler TTVs

### v1

@@ 3.3 Radial-velocity modeling
After revealing planet-induced Doppler signals with periodogram analyses of the HARPS-N RV dataset of Kepler-10, we fitted them with a multi-Keplerian model by i) varying the number of planets in the system from two to four; ii) employing three different Bayesian modeling approaches (see Sects. 3.3.1-3.3.3) in order to evaluate the robustness of the orbital solution, given the previous discrepancies in the literature; iii) using a Gaussian likelihood function, which accounts for additional uncorrelated (white) noise (e.g., Gregory 2005) and/or correlated (red) noise through Gaussian Process (GP) regression with quasi-periodic and/or squared exponential covariance functions (e.g., Grunblatt et al. 2015; Rajpaul et al. 2015).
@@
The Keplerian signal of each planet was modeled with five parameters: the orbital period $P$; the inferior conjunction time $T_{\rm c}$, which is equivalent to the mid-transit time for the transiting planets; $\sqrt{e}\cos(\omega)$ and $\sqrt{e}\sin(\omega)$, where $e$ and $\omega$ are the orbital eccentricity and the argument of periastron; and the RV semi-amplitude $K$. A zero point $\gamma$ and a jitter term $\sigma_{\rm j}$ were also fitted for each dataset. The jitter terms were summed in quadrature to the formal RV uncertainties to take additional uncorrelated noise of unknown origin (stellar, instrumental, etc.) into account.
@@
To check whether possible correlated noise is also present in the data and might be modeled along with the planetary signals, despite the very low stellar magnetic activity level, we further considered Gaussian Processes with the following hyper-parameters: the amplitude $h$ of the GP, the exponential decay term $\lambda$, the periodic GP term $P_{\rm rot}$, and the inverse harmonic complexity term $w$. The first two hyper-parameters are in common between the quasi-periodic (QP) and squared exponential (SE) covariance functions (kernels).
@@
  • half-Gaussian priors centered on zero with $\sigma_{\rm e}=0.098$ for the orbital eccentricities of Kepler-10 c and Kepler-10 d, following the distribution of eccentricities in multi-planet systems (Van Eylen et al. 2019). This prior avoids spurious high eccentricities, and hence dynamical instabilities due to orbit crossings; spurious eccentricities may occur when the RV semi-amplitudes are comparable to the noise level as in the present case (e.g., Zakamska et al. 2011; Hara et al. 2019).
    For the innermost planet, Kepler-10 b, we used instead a circular orbit, because its orbit is expected to be well circularized, given its extremely small (short) semi-major axis (orbital period) (e.g., Matsumura et al. 2008, 2010).
@@ 3.3.1 Differential evolution Markov chain Monte Carlo (DE-MCMC)
The DE-MCMC method is the Markov chain Monte Carlo (MCMC) version of the differential evolution (DE) genetic algorithm (Ter Braak 2006; Eastman et al. 2013, 2019). A number of chains equal to twice the number of free parameters are run simultaneously, and perform an optimized exploration of the parameter space through the automatic choice of step scales and orientations. The sampling distribution for each step of a given chain consists in i) randomly selecting two other chains; ii) computing the difference of the model parameters of these two chains; iii) determining the proposal step from ii) and Eq. 2 in Ter Braak (2006); iv) accepting or rejecting the proposal step according to the Metropolis-Hastings algorithm. For the identification of the burn-in steps and the convergence and well mixing of the DE-MCMC chains we followed the criteria proposed by Eastman et al. (2013).
@@
We ran a DE-MCMC analysis first on the HARPS-N RVs only, and then on the combined HARPS-N and HIRES datasets. We accounted for uncorrelated (white) noise only, because the GP analysis with either QP or SE kernels did not achieve convergence of the chains, but led to very low acceptance rates and unconstrained hyper-parameters, as expected from the low level of stellar magnetic activity.
@@
The results of the DE-MCMC for the various planet models are listed in Table LABEL:table_harpsn_param for the HARPS-N RVs. We used the Bayesian Information Criterion (BIC) (e.g., Kass & Raftery 1995; Jeffreys 1998; Mukherjee et al. 1998; Burnham & Anderson 2004; Liddle 2007) to compute the relative probabilities of the models with two, three and four planets, and found $BIC_{\rm 3pl}-BIC_{\rm 2pl}=-10.2$ and $BIC_{\rm 4pl}-BIC_{\rm 3pl}=+3.3$. This implies that the 3-planet model is favored over both the 2-planet and 4-planet models, being $\sim 165$ and 5 times more likely, respectively. Moreover, an additional planet with $P\sim 83$ d in the 4-planet model would not be compatible with the TTVs of Kepler-10 c (see Sect. 3.5). A very important outcome is that the values of the RV semi-amplitudes $K$ of planet b and, especially, c do not vary significantly with the adopted (2- or 3-planet) model, being fully consistent within $1\sigma$ (see Table LABEL:table_harpsn_param).
@@
Adding the C2-decker HIRES data does not change significantly the planet parameters, as expected from their low number compared to the HARPS-N RV measurements (47 vs 291). The RV semi-amplitudes agree within $1\sigma$ with the values obtained with HARPS-N only (see Table LABEL:table_harpsn-hires_param), though the RV semi-amplitudes of both Kepler-10 c and Kepler-10 d get slightly lower by $1-\sigma$ and have slightly larger relative errors; this might be due to some HIRES instrumental effects, the investigation of which goes beyond the scope of this work. In any case, we mainly rely on the solution obtained from the modeling of the HARPS-N data, given that the correction of systematics and/or low-level activity variations with YARARA-v2 could be applied to the HARPS-N spectra only.
@@
The best-fit with the 3-planet model from the DE-MCMC analysis is shown in Fig. 6 for the HARPS-N RVs, and in Fig. 17 for the HARPS-N and HIRES data.
@@ 3.3.2 Nested Sampling through MultiNest
The Kepler-10 RVs were also modeled using the publicly available Monte Carlo nested sampler and Bayesian inference tool MultiNest v3.10 (e.g., Feroz et al. 2019), through the pyMultiNest wrapper (Buchner et al. 2014). The set-up was characterised by 500 live points, a sampling efficiency of 0.3, and a Bayesian tolerance of 0.5. The Bayesian model selection was performed by comparing the values of the Bayesian evidence $\ln\mathcal{Z}$ calculated by MultiNest by assigning the same a-priori probability to each model.
We found that i) the 3-planet model is strongly favored over the 2-planet model with $\ln\mathcal{Z}_{3\,pl.}-\ln\mathcal{Z}_{2\,pl.}$=8.4, and ii) the 4-planet model is weakly favored over the 3-planet model, given $\ln\mathcal{Z}_{4\,pl.}-\ln\mathcal{Z}_{3\,pl.}$=1.1 and the empirical scale reported in Feroz et al. (2011); the posterior distribution of the fourth signal converged to a period of $P=82.95^{+0.20}_{-4.64}$ d, encompassing the 1-yr alias of the main mode at $83$ d.
3.3.3 Nested sampling through PolyChord with and without multi-dimensional Gaussian Processes
The Kepler-10 RVs show only weak (linear) correlation with activity indicators, as expected for an old, quiet star. Specifically, the Pearson’s coefficient between RV and BIS ($\log R^{\prime}_{\rm HK}$) measurements is $r=23.1\%,p=0.09$ ($r=-8.5\%,p=0.55$) before the CCD replacement, and $r=-15.0\%,p=0.021$ ($r=4.0\%,p=0.54$) afterwards. Nevertheless, we used the multi-dimensional (MD) GP framework developed by Rajpaul et al. (2015) to model RVs simultaneously with ${\rm BIS}$ and $\log R^{\prime}_{\rm HK}$ observations. This GP framework assumes that any observed stellar activity signals are generated by some underlying latent function and its derivatives; this function, not observed directly, is modelled with a GP (Rasmussen & Williams 2006; Roberts et al. 2013). The framework includes components to account for convective blueshift suppression and active region evolution.
@@ 3.3.3 Nested sampling through PolyChord with and without multi-dimensional Gaussian Processes
For the computation of the posterior probabilities of the model parameters we used PolyChord, a nested sampling algorithm designed to work well even with parameter spaces with very large dimensionality (Handley et al. 2015). For a short discussion illuminating its favourable properties compared with MultiNest, its direct predecessor, see Hall et al. (2018); a detailed study testing PolyChord and validating its use for joint modelling of exoplanets and stellar activity in RVs is given by Ahrer et al. (2021). We called PolyChord via the pypolychord Python wrapper, leaving sampling parameters at their default values, except for the stopping (precision) criterion, which we changed from the default $10^{-3}$ to a (much) more stringent $10^{-9}$, to minimise the scatter in model evidence values from run to run (Ahrer et al. 2021). We ran all PolyChord sampling on a high-performance computing platform, typically using several hundred CPU cores simultaneously.
@@
In addition to the MD GP analysis, we also used PolyChord with models describing RVs only, without any GP component: partly as a comparison with the multi-dimensional GP analysis, and partly to be able to cross-check results obtained with other approaches (e.g., MultiNest, Sect. 3.3.2). We applied both classes of models (multi-dimensional GP; no GP) to HARPS-N data only, and to HARPS-N plus HIRES data, for a total of four separate modelling ‘setups’, (i)–(iv). For each of these four setups, we explored models with between 00 and $4$ Keplerian terms (each model assumed equally likely a priori), and performed three separate PolyChord runs for each model within a setup to obtain robust constraints on evidences. The results from this analysis appear in Table 11.
@@
Across the four setups, planets b and c were reassuringly and resoundingly detected ($\Delta\ln\mathcal{Z}\sim 30$, compared to models without either planet). A model incorporating three Keplerians (planets b and c, plus one non-transiting planet) was strongly favoured over 2-planet models across all four setups, with $\Delta\ln\mathcal{Z}\sim 6$. As for planet orbital parameters, $K_{b}$ was strongly consistent while $K_{c}$ was broadly consistent (within about $1\sigma$) across all setups and applicable models. For the third Keplerian, we obtained a period of $P_{\rm d}\sim 151$ d (with a small secondary peak at $124.5$ d) and semi-amplitude $K_{\rm d}\sim 1.4$ $\textrm{m}\leavevmode\nobreak\ \textrm{s}^{-1}$ across all setups and applicable models – this consistency lent additional weight to the conclusion that the third Keplerian term corresponds to a genuine planet. By contrast, there was insufficient statistical evidence for a fourth planet in any of our setups, with $\Delta\ln\mathcal{Z}\lesssim 1$ for the best $4$-planet versus the best $3$-planet models. For a fourth Keplerian, the period posterior distributions peaked at $P\sim$83 d (corresponding semi-amplitude $K\sim 0.8$ $\textrm{m}\leavevmode\nobreak\ \textrm{s}^{-1}$) with a prominent secondary peak at $114$ d.
@@
Given the strong consistency between the results we obtained with the different modelling setups, and for direct comparison with results from other methods, we present in Table LABEL:table_harpsn_param the results from the multi-dimensional GP analysis applied to HARPS-N data only. We note that our planet parameter posteriors were generally broader when including HIRES data, with evidences in favour of the detection of the known planets b and c weakened slightly (mirroring Rajpaul et al. 2021). We also note that the multi-dimensional GP modeling with HARPS-N data only led us to a GP period of $54^{+10}_{-9}$ d, and a $1\sigma$ upper limit on the activity RV semi-amplitude of $<1.32$ $\textrm{m}\leavevmode\nobreak\ \textrm{s}^{-1}$. As already pointed out by Rajpaul et al. (2017), it is nevertheless unclear whether the GP period of $\sim 54$ d corresponds to the stellar rotation period, given that neither the Kepler light curve nor the time series of the CCF and CaII activity indicators show clear signals attributable to the stellar rotation.
@@ 3.4 Search for Kepler-10 d transits
Given the presence of a significant signal at $\sim 151$ d, we analyzed the Kepler light curve to check whether any transit could have been missed in previous studies of this system. In particular, we first modeled the light curves using the python package wotan (Hippke et al. 2019) and then performed a Transit Least Squares (TLS) periodogram (Hippke & Heller 2019) on the residuals of the light curve, after modeling the transits of both Kepler-10 b and Kepler-10 c. We found no significant periodicity at $\sim 151$ d or elsewhere, despite the Kepler curve being about $\sim 1480$ d long (minus a few gaps in-between). We also visually inspected the light curve around the 8 predicted transits of Kepler-10 d (from the ephemeris of Table LABEL:table_harpsn_param) and found no significant dip. We point out that, based on the estimated RV semi-amplitude of the outer planet, we would expect transits of comparable depth to planet c ($\sim 0.5$ mmag), and thus easily detectable (see Figure 7). Planet d should transit if it had the same inclination as planet c, since $b_{d}\leq 1$:
@@
Figure 7: Kepler light curve phase-folded to the period of planet d (from Table LABEL:table_harpsn_param). No transit-like features with a depth comparable to planet c (blue horizontal line) can be appreciated.

### v2

@@ 3.4 Modeling of radial velocities
After revealing planet-induced Doppler signals with periodogram analyses of the HARPS-N RV dataset of Kepler-10, we fitted them with a multi-Keplerian model by (i) varying the number of planets in the system from two to four; (ii) employing three different Bayesian modeling approaches (see Sects. 3.4.1-3.4.3) in order to evaluate the robustness of the orbital solution, given the previous discrepancies in the literature; and (iii) using a Gaussian likelihood function, which accounts for additional uncorrelated (white) noise (e.g., Gregory 2005) and/or correlated (red) noise through GP regression with quasi-periodic (QP) and/or squared exponential (SE) covariance functions (e.g., Grunblatt et al. 2015; Rajpaul et al. 2015).
@@
The Keplerian signal of each planet was modeled with five parameters: the orbital period $P$; the inferior conjunction time $T_{\rm c}$, which is equivalent to the mid-transit time for the transiting planets; $\sqrt{e}\cos(\omega)$ and $\sqrt{e}\sin(\omega)$, where $e$ and $\omega$ are the orbital eccentricity and the argument of periastron; and the RV semiamplitude $K$. A zero point $\gamma$ and a jitter term $\sigma_{\rm j}$ were also fitted for each dataset. The jitter terms were summed in quadrature to the formal RV uncertainties to take additional uncorrelated noise of unknown origin (stellar, instrumental, etc.) into account.
@@
To check whether possible correlated noise is also present in the data and might be modeled along with the planetary signals, despite the very low stellar magnetic activity level, we further considered Gaussian Processes with the following hyperparameters: the amplitude $h$ of the GP, the exponential decay term $\lambda$, the periodic GP term $P_{\rm rot}$, and the inverse harmonic complexity term $w$. The first two hyperparameters are in common between the QP and SE covariance functions (kernels).
@@
  • half-Gaussian priors centered on zero with $\sigma_{\rm e}=0.098$ for the orbital eccentricities of Kepler-10 c and Kepler-10 d, following the distribution of eccentricities in multiplanet systems (Van Eylen et al. 2019). This prior avoids spurious high eccentricities, and hence dynamical instabilities due to orbit crossings; spurious eccentricities may occur when the RV semiamplitudes are comparable to the noise level as in the present case (e.g., Zakamska et al. 2011; Hara et al. 2019).
    For the innermost planet, Kepler-10 b, we used instead a circular orbit, because its orbit is expected to be well circularized given its extremely small (short) semimajor axis (orbital period; e.g., Matsumura et al. 2008, 2010).
@@ 3.4.1 Differential evolution Markov chain Monte Carlo (DE-MCMC)
The DE-MCMC method is the Markov chain Monte Carlo (MCMC) version of the differential evolution (DE) genetic algorithm (Ter Braak 2006; Eastman et al. 2013, 2019). A number of chains equal to twice the number of free parameters are run simultaneously, and perform an optimized exploration of the parameter space through the automatic choice of step scales and orientations. The sampling distribution for each step of a given chain consists in (i) randomly selecting two other chains; (ii) computing the difference of the model parameters of these two chains; (iii) determining the proposal step from (ii) and Eq. 2 in Ter Braak (2006); and (iv) accepting or rejecting the proposal step according to the Metropolis-Hastings algorithm. For the identification of the burn-in steps and the convergence and well mixing of the DE-MCMC chains, we followed the criteria proposed by Eastman et al. (2013).
@@
We ran a DE-MCMC analysis first on the HARPS-N RVs only, and then on the combined HARPS-N and HIRES datasets. We accounted for uncorrelated (white) noise only, because the GP analysis with either QP or SE kernels did not achieve convergence of the chains, but led to very low acceptance rates and unconstrained hyperparameters, as expected from the low level of stellar magnetic activity.
@@
The results of the DE-MCMC for the various planet models are listed in Table LABEL:table_harpsn_param for the HARPS-N RVs. We used the Bayesian information criterion (BIC; e.g., Kass & Raftery 1995; Jeffreys 1998; Mukherjee et al. 1998; Burnham & Anderson 2004; Liddle 2007) to compute the relative probabilities of the models with two, three, and four planets, and found $BIC_{\rm 3pl}-BIC_{\rm 2pl}=-10.2$ and $BIC_{\rm 4pl}-BIC_{\rm 3pl}=+3.3$. This implies that the three-planet model is favored over both the two-planet and four-planet models, being $\sim 165$ and 5 times more likely, respectively. Moreover, an additional planet with $P\sim 83$ d in the four-planet model would not be compatible with the TTVs of Kepler-10 c (see Sect. 3.6). A very important outcome is that the values of the RV semiamplitudes $K$ of planet b and, especially, c do not vary significantly with the adopted (two- or three-planet) model, being fully consistent within $1\sigma$ (see Table LABEL:table_harpsn_param).
@@
Adding the C2-decker HIRES data does not change significantly the planet parameters, as expected from their low number compared to the HARPS-N RV measurements (47 vs 291). The RV semiamplitudes agree within $1\sigma$ with the values obtained with HARPS-N only (see Table LABEL:table_harpsn-hires_param), though the RV semiamplitudes of both Kepler-10 c and Kepler-10 d get slightly lower by $1\sigma$ and have slightly larger relative errors; this might be due to some HIRES instrumental effects, the investigation of which goes beyond the scope of this work. In any case, we mainly rely on the solution obtained from the modeling of the HARPS-N data, given that the correction of systematics and/or low-level activity variations with YARARA-v2 could be applied to the HARPS-N spectra only.
@@
The best-fit with the three-planet model from the DE-MCMC analysis is shown in Fig. 6 for the HARPS-N RVs, and in Fig. 16 for the HARPS-N and HIRES data.
@@ 3.4.2 Nested sampling through MultiNest
The Kepler-10 RVs were also modeled using the publicly available Monte Carlo nested sampler and Bayesian inference tool MultiNest v3.10 (e.g., Feroz et al. 2019), through the pyMultiNest wrapper (Buchner et al. 2014). The setup was characterized by 500 live points, a sampling efficiency of 0.3, and a Bayesian tolerance of 0.5. The Bayesian model selection was performed by comparing the values of the Bayesian evidence $\ln\mathcal{Z}$ calculated by MultiNest by assigning the same a-priori probability to each model.
We found that (i) the three-planet model is strongly favored over the two-planet model with $\ln\mathcal{Z}_{3\,pl.}-\ln\mathcal{Z}_{2\,pl.}$=8.4, and (ii) the four-planet model is weakly favored over the three-planet model, given $\ln\mathcal{Z}_{4\,pl.}-\ln\mathcal{Z}_{3\,pl.}$=1.1 and the empirical scale reported in Feroz et al. (2011); the posterior distribution of the fourth signal converged to a period of $P=82.95^{+0.20}_{-4.64}$ d, encompassing the 1-yr alias of the main mode at $83$ d.
3.4.3 Nested sampling through PolyChord with and without multidimensional Gaussian processes
The Kepler-10 RVs show only weak (linear) correlation with activity indicators, as expected for an old, quiet star. Specifically, the Pearson’s coefficient between RV and BIS ($\log R^{\prime}_{\rm HK}$) measurements is $r=23.1\%,p=0.09$ ($r=-8.5\%,p=0.55$) before the CCD replacement, and $r=-15.0\%,p=0.021$ ($r=4.0\%,p=0.54$) afterwards. Nevertheless, we used the multidimensional (MD) GP framework developed by Rajpaul et al. (2015) to model RVs simultaneously with ${\rm BIS}$ and $\log R^{\prime}_{\rm HK}$ observations. This GP framework assumes that any observed stellar activity signals are generated by some underlying latent function and its derivatives; this function, not observed directly, is modeled with a GP (Rasmussen & Williams 2006; Roberts et al. 2013). The framework includes components to account for convective blueshift suppression and active region evolution.
@@ 3.4.3 Nested sampling through PolyChord with and without multidimensional Gaussian processes
For the computation of the posterior probabilities of the model parameters, we used PolyChord, a nested sampling algorithm designed to work well even with parameter spaces with very large dimensionality (Handley et al. 2015). A short discussion illuminating its favorable properties compared with MultiNest, its direct predecessor, can be found in Hall et al. (2018); a detailed study testing PolyChord and validating its use for joint modeling of exoplanets and stellar activity in RVs is given by Ahrer et al. (2021). We called PolyChord via the pypolychord Python wrapper, leaving sampling parameters at their default values, except for the stopping (precision) criterion, which we changed from the default $10^{-3}$ to a (much) more stringent $10^{-9}$, to minimize the scatter in model evidence values from run to run (Ahrer et al. 2021). We ran all PolyChord sampling on a high-performance computing platform, typically using several hundred CPU cores simultaneously.
@@
In addition to the MD GP analysis, we also used PolyChord with models that describe RVs only, without any GP component: partly as a comparison with the MD GP analysis, and partly to be able to crosscheck results obtained with other approaches (e.g., MultiNest, Sect. 3.4.2). We applied both classes of models (MD GP; no GP) to HARPS-N data only, and to HARPS-N plus HIRES data, for a total of four separate modeling ‘setups’, (i)–(iv). For each of these four setups, we explored models with between 00 and $4$ Keplerian terms (each model assumed equally likely a priori), and performed three separate PolyChord runs for each model within a setup to obtain robust constraints on the evidences. The results from this analysis appear in Table 11.
@@
Across the four setups, planets b and c were reassuringly and resoundingly detected ($\Delta\ln\mathcal{Z}\sim 30$, compared to models without either planet). A model incorporating three Keplerians (planets b and c, plus one non-transiting planet) was strongly favored over two-planet models across all four setups, with $\Delta\ln\mathcal{Z}\sim 6$. As for the planet orbital parameters, $K_{b}$ was strongly consistent while $K_{c}$ was broadly consistent (within about $1\sigma$) across all setups and applicable models. For the third Keplerian, we obtained a period of $P_{\rm d}\sim 151$ d (with a small secondary peak at $124.5$ d) and semiamplitude $K_{\rm d}\sim 1.4$ $\textrm{m}\leavevmode\nobreak\ \textrm{s}^{-1}$ across all setups and applicable models – this consistency lent additional weight to the conclusion that the third Keplerian term corresponds to a genuine planet. By contrast, there was insufficient statistical evidence for a fourth planet in any of our setups, with $\Delta\ln\mathcal{Z}\lesssim 1$ for the best $4$-four-planet versus the best three-planet models. For a fourth Keplerian, the period posterior distributions peaked at $P\sim$83 d (corresponding semiamplitude $K\sim 0.8$ $\textrm{m}\leavevmode\nobreak\ \textrm{s}^{-1}$) with a prominent secondary peak at $114$ d.
@@
Given the strong consistency between the results we obtained with the different modeling setups, and for direct comparison with results from other methods, we present in Table LABEL:table_harpsn_param the results from the MD GP analysis applied to HARPS-N data only. We note that our planet parameter posteriors were generally broader when including HIRES data, with evidence in favor of the detection of the known planets b and c weakened slightly (mirroring Rajpaul et al. 2021). We also note that the MD GP modeling with HARPS-N data only led us to a GP period of $54^{+10}_{-9}$ d, and a $1\sigma$ upper limit on the activity RV semiamplitude of $<1.32$ $\textrm{m}\leavevmode\nobreak\ \textrm{s}^{-1}$. As already pointed out by Rajpaul et al. (2017), it is nevertheless unclear whether the GP period of $\sim 54$ d corresponds to the stellar rotation period, given that neither the Kepler light curve nor the time series of the CCF and CaII activity indicators show clear signals attributable to the stellar rotation.
@@ 3.5 Search for Kepler-10 d transits
Given the presence of a significant signal at $\sim 151$ d, we analyzed the Kepler light curve to check whether any transit could have been missed in previous studies of this system. In particular, we first modeled the light curves using the python package wotan (Hippke et al. 2019) and then performed a transit least squares (TLS) periodogram (Hippke & Heller 2019) on the residuals of the light curve, after modeling the transits of both Kepler-10 b and Kepler-10 c. We found no significant periodicity at $\sim 151$ d or elsewhere, despite the Kepler curve being about $\sim 1480$ d long (minus a few gaps in between). We also visually inspected the light curve around the eight predicted transits of Kepler-10 d (from the ephemeris of Table LABEL:table_harpsn_param) and found no significant dip. We point out that, based on the estimated RV semiamplitude of the outer planet, we would expect transits of a comparable depth to planet c ($\sim 0.5$ mmag), and thus easily detectable (see Fig. 7). Planet d would transit if it had the same inclination as planet c, since $b_{d}\leq 1$:
@@
Figure 7: Kepler light curve phase-folded to the period of planet d (from Table LABEL:table_harpsn_param). No transit-like features with a depth comparable to planet c (blue horizontal line) can be seen.